CLEAR Weekly Operation Meeting
                    
    
        
            
            →
            
        
        Europe/Zurich
    
                
                
                    
                        
                            
    
    
        
            
                
                CERN
            
            
                
    
        
        
            
        
    
                        
                    
                
                
                
            CERN
MINUTES
Actions from previous week:
- Check the availability of a beam pipe to allow the installation of only two CLIC cavity BPM > ongoing
 - Further look to solve the noise in the plasma lens pump.
 - Devise a way of doing laser alignment on the in-air test stand.
 - Understand where's the end of the trigger cables in the klystron gallery.
 - Look at moving one of the unused CTF3 front ends to the barrack to aid data collection.
 
New actions:
- Reconnection of the thermalized water pipes on the X-band ACS
 - Re-installation of safety valve on gas pipe for P.L.
 
02/05/2021, 18:52 Roberto Corsini
                    There are minutes attached to this event.
                    Show them.